# Teuken-7B-Base & Teuken-7B-Instruct: Towards European LLMs

### v2

@@ Abstract
We present two multilingual Large Language Models designed to embrace Europe’s linguistic diversity by supporting all 24 official languages of the European Union. Trained on a dataset comprising around 60% non-English data and utilizing a custom multilingual tokenizer, our models address the limitations of existing LLMs that predominantly focus on English or a few high-resource languages. We detail the models’ development principles, i.e., data composition, tokenizer optimization, and training methodologies. The models demonstrate competitive performance across multilingual benchmarks, as evidenced by their performance on European versions of ARC, HellaSwag, MMLU, and TruthfulQA.
Teuken-7B-Base & Teuken-7B-Instruct: Towards European LLMs
Mehdi Ali1,2, Michael Fromm1,2, Klaudia Thellmann3, Jan Ebert4, Alexander Arno Weber1,2, Richard Rutmann1,2, Charvi Jain1,2, Max Lübbering1,2, Daniel Steinigen1, Johannes Leveling1, Katrin Klug1, Jasper Schulze Buschhoff1, Lena Jurkschat3, Hammam Abdelwahab1, Benny Jörg Stein1, Karl-Heinz Sylla1, Pavel Denisov1, Nicolo’ Brandizzi1, Qasid Saleem1, Anirban Bhowmick1, Lennard Helmer1, Chelsea John4, Pedro Ortiz Suarez5, Malte Ostendorff5, Alex Jude1, Lalith Manjunath3, Samuel Weinbach7, Carolin Penke4, Oleg Filatov4, Shima Asaadi6, Fabio Barth5, Rafet Sifa1, Fabian Küch6, Andreas Herten4, René Jäkel3, Georg Rehm5, Stefan Kesselheim4, Joachim Köhler1, Nicolas Flores-Herr1
1Fraunhofer IAIS, 2Lamarr Institute, 3TU Dresden,4FZ Jülich, 5DFKI, 6Fraunhofer IIS, 7Aleph Alpha
thanks: †Equal contribution.
@@ 1 Introduction
LLMs represents a disruptive technology that has the potential to be applied in numerous applications. To develop LLMs, expertise in various areas is required, starting from large-scale data pre-processing Penedo et al. (2024), training efficient tokenizers Petrov et al. (2023); Ali et al. (2024), pre-training the models efficiently on a vast infrastructure spanning thousands of GPUs, instruction tuning the pre-trained models, and properly evaluating them on various downstream tasks Dubey et al. (2024). Multilingualism as an additional dimension introduces additional considerations to all phases of the model development, such as multilingual data composition and multilingual tokenizer. Therefore, it is crucial that the technology and expertise to build these models is democratized to enable different communities and organizations to employ these models for their use cases.
Many efforts in developing open-source LLMs have been undertaken, such as BLOOM Scao et al. (2022), LLaMA-3 Dubey et al. (2024), OLMo Groeneveld et al. (2024), Aya Üstün et al. (2024), and Mistral Jiang et al. (2023).
While these existing efforts represent significant contributions to the community and the advancement of artificial intelligence, there are still two major limitations. First, the current open-source models are predominantly English-centric, limiting their use to a broad multilingual context covering high and low resource languages such as within the European Union. Relying on English-centric models that employ an English-centric tokenizer in a multilingual context introduces severe disadvantages such as lower downstream performance, additional inference costs, and increased training costs for language-specific continued pre-training and fine-tuning for languages besides English Petrov et al. (2023); Ali et al. (2024). Second, the open-source efforts disclose different levels of granularity regarding sharing details about the development of the models. Even though information regarding the model architecture is usually shared, the decisions/ablation experiments behind certain architectural choices are not always described, which can provide crucial insights relevant to developing custom LLMs. This limitation is even more apparent in the description of the dataset composition, which often provides a coarse overview, hampering the reproduction of the work.
@@
  • We present Teuken-7B-Base, a European pre-trained LLM that has been trained from scratch for 4 trillion tokens based on all 24 official European languages.
  • We present Teuken-7B-Instruct, the is instruction tuned model on top of the base model
Besides these artefacts, we describe our design decisions in detail to facilitate the reproduction of our work and the development of novel models based on our work. Therefore, our contributions present an essential step towards the democratization of technology within Europe.
@@ 2 Related Work
Since the introduction of GPT-3 Brown et al. (2020), several open-source/open-weights efforts have been undertaken to train LLMs. While the large majority of the work focus on English-centric models Zhang et al. (2022); Touvron et al. (2023a, b); Groeneveld et al. (2024); Dubey et al. (2024), there have been also efforts training multilingual models.
One of the most prominent examples is BLOOM Scao et al. (2022), a 176B LLM trained on 46 natural languages. Further examples that specifically address multilingualism are the encoder-decoder models mT5 Xue et al. (2021), XLM Lample and Conneau (2019), XLM-R Conneau et al. (2020), and the encoder model mBERT Devlin et al. (2019).
Unlike the previously mentioned efforts, we specifically address 24 official European languages and focus on ensuring that a large fraction of the training data is composed of non-English data, representing a major step towards European LLMs. Concurrent to our work, EuroLLM Martins et al. (2024), a 1.7B decoder-only LLM that follows the same spirit as our undertaking by addressing all 24 European languages and Salamandra [^1] https://huggingface.co/BSC-LT/salamandra-7b covering 32 languages, have been presented.
@@ 3 Pre-Training Data
We used the dataset presented in Brandizzi et al. (2024) as a base since it i.) covers all 24 official EU languages, ii.) is large, iii.) contains a large amount of non-English data, iv.) comprises various domains, and v.) is filtered based on established practices. In the following, we describe the composition of our final training comprised of web-crawled and curated data.
For the the web crawled part, we sampled our dataset based on the 60 filtered Common Crawls WET dumps (cf.  Section A.1) provided by Brandizzi et al. Brandizzi et al. (2024). The raw WET dumps have been filtered based on established heuristics, ensuring that the pre-training corpus is of high quality Brandizzi et al. (2024). Additionally, we employed the recently released Fine-Web EDU dataset Penedo et al. (2024) as significant performance improvements for models trained on the dataset have been reported. It should be noted that Fine-Web EDU is only available in English. We still integrated it and investigated the cross-lingual performance of the corresponding model checkpoint (see Section 5.2). To ensure a multilingual composition, we up-sampled all languages except English, which we down-sampled. Specifically, we employed the content for all languages except English Overall, the web crawled part makes up 86.55% (cf.  Table 42) our pre-training dataset.
In addition to web crawled data, we included all curated datasets from Brandizzi et al. (2024). Part of the curated datasets are domain-specific data such as academics, finance, and patents ensuring the diversity of our training dataset.
Our composed training dataset contains 4 trillion tokens, of which 86.55% (cf. Table 42) originates from web data, and the remaining 13.45% represent is curated data As illustrated in Figure 1 and Figure 2, 41.70% of the tokens stem from English content and due to the inclusion of German, French, and Spanish, we approach around two-thirds of the total tokens.
@@ 4 Multilingual Tokenization and Fertility Impact on Model Efficiency
In multilingual natural language processing (NLP), it is crucial to train balanced multilingual tokenizers Petrov et al. (2023); Ali et al. (2024) to avoid increased training and inference costs and latency during inference for non-English queries. Furthermore, it prevents the model from learning long-range dependencies in limited context windows Vaswani et al. (2017).
Therefore, we developed a custom multilingual tokenizer, closely following Ali et al. (2024), that is optimized for all 24 official European languages. It aims to reduce excessive text fragmentation, a phenomenon termed high “fertility”, and refers to the average number of tokens generated per word.
Fertility ($F$) is defined as the ratio of the total number of tokens ($T$) to the total number of words ($W$) in a text, as shown in the following:
$F=\frac{T}{W}$
We conducted a fertility analysis on 2,000 sentences from the FLORES-200 Costa-jussà et al. (2022) dataset to compare tokenizers. Because the dataset is translated across languages, i.e., the analysis is conducted on semantic equivalent content, it provides a reliable basis for evaluation. A comparison with other widely-used tokenizers is presented in Figure 3
@@ 5 Base Model
In the following, we describe the model architecture (Section 5.1) and training (Section 5.2). Additionally, we describe in the appendix the used training framework (Section A.7) and the training infrastructure (Section A.8).
@@ 5.1 Model Architecture
Our model is a 7B transformer-based decoder-only model. Table 7 provides an overview of our model architecture. We want to highlight that our architectural choices are derived from internal ablation studies and findings from related work. Our models have a sequence length of 4096 tokens and employ Rotary Su et al. (2024) positional embeddings that are employed to train state-of-the-art models Dubey et al. (2024). To accelerate inference and reduce memory requirements, we employed grouped-query attention Ainslie et al. (2023). An entire overview of our architectural choices is presented in Table 7.
@@
These and other design decisions were guided by medium-scale (Chinchilla-optimal Hoffmann et al. (2022) training of a 2.6B parameter model) ablation runs for various training-related hyperparameters. Our goal with these ablations was to find improvements in the compute-equivalent setting while also confirming whether proposed modifications transfer to our codebase, which is not necessarily the case Narang et al. (2021). Appendix A.2 describes the conducted ablation experiments in detail, Table 6 summarizes the results, and Figures 6 and 7 contain loss curves for the various experiments.
5.2 Training
@@
Using the causal language modelling training objective, we trained our model on 4T tokens covering all 24 European languages as described in Section 3, employed AdamW as an optimizer and used a cosine learning rate schedule starting with a learning rate of 3e-5, increasing it to the maximum learning rate of 3e-4 within the first 10,000 steps, and decaying it afterwards. During the training, we took two additional design decisions motivated by recent findings from related work.
@@
First, Penedo et al. Penedo et al. (2024) showed that training based on educational content further improves the performance of LLMs. Therefore, after 2.85T tokens, we conducted an ablation where we trained one model up to 3T tokens based on FineWeb-EDU Penedo et al. (2024) (which is only available in English), and a second model with the initial data composition (see Section 3). As shown in Figure 4, the model trained based on education content obtained a performance % across all tasks and languages except for HellaSwag, where the results were comparable. In the appendix, we show the performance development for English (Figure 8), French (Figure 9), German (Figure 10), Finish (Figure 11) and Estonian (Figure 12).
Second, we want to emphasize that the initial goal has been to train our base model up to 3T tokens, which are beyond compute-optimal based on the scaling laws presented by Hoffmann et al. Hoffmann et al. (2022) for our model size. Because recent works show that it is beneficial to further the pre-train model, e.g., LLaMA 3 has been trained up to 15T tokens, we decided to train the model on an additional 1T token. Consequently, we faced the challenge of defining the corresponding learning rate schedule that had been originally defined for training up to 3T tokens. We ablated different settings to continue pre-training inspired by Ibrahim et al. (2024) as described in Section A.4. As a result, we further decreased the learning rate from 10% of the maximum learning rate to 8.3% within the following 1T tokens.
@@
Figure 4: Downstream performance of the base model from 0.99T to 4T tokens across 21 European languages. The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens.
@@ 6 Instruction Tuned Model
In this section, we describe the creation of the instruction tuned model, including the dataset composition (Section 6.1) and the instruction tuning (Section 6.2). Additionally to "Ours (Instruct)", we ablated another model variant, which contains examples for all 24EU Languages, as shown in Table 41. To our surprise and as visible in  Figure 14, both variants are comparable in performance on MT-Bench-X. However, "Ours (Instruct)" performs slightly better on the EU21 benchmark for selected languages. Hence, we present the variant "Ours (Instruct)", which focuses on English and German for instruction-tuning.
@@ 6.1 Data
In the work of Weber et al. (2024), the authors show that employing (parallel) multilingual instruction tuning datasets improves cross-lingual instruction following the capabilities of the model. Therefore, we composed a multilingual dataset of English German data. The dataset comprises several subsets, as listed in Table 40. Our dataset is not entirely parallel, as suggested by  Weber et al. (2024) due to the lack of parallel data.
English
Our objective was to balance the dataset, aiming to include a similar number of English and German examples. As shown with the subset selection in Table 40, we included only high-quality English data. To reduce the dataset size, the following multi-step filtering process was executed: First, we computed reward scores for all English examples using Starling-RM-7B-alpha [^2] https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha. Second, to increase the models’ coding ability and multi-turn chat capability, we include all multi-turn examples, the complete code alpaca‘ subset, and the portion of the LMSYS-Chat dataset subset noted as high-quality by  Zheng et al. (2024). The remaining high-quality subsets — Open Orca (Cf. Mukherjee et al. (2023)) and both versions of the Evol instruct dataset Xu et al. (2024) — are then sampled to balance the German and English dataset proportion. For this, we select the highest-scoring examples to ensure that each subset contributed an equal amount of high-quality data.
German
For the German data selection, we incorporated several datasets to ensure broad linguistic coverage. We included the German and English subsets of Bactrian-X dataset, as Weber et al. (2024) demonstrated that semantically parallel data enhances cross-lingual model performance. In the same spirit, we include a proportion of the Alpaca, Evol-Instruct, and ShareGPT datasets for which either the instructions or also the responses are translated to German by Chen et al. (2023).
6.2 Post-Training
For instruction-tuning the model, we extended FastChat [^3] https://github.com/lm-sys/FastChat/tree/main/fastchat for multilingual system prompts and translated the system prompt “A chat between a human and an artificial intelligence assistant. The assistant gives helpful and polite answers to the human’s questions.” for all 24 official EU languages. We employed standard instruction tuning by calculating the loss only for the assistant responses of a multi-turn conversation. We trained the model on 8xH100 GPUs for 2.5 days for three epochs.

### v3

@@ Abstract
We present two multilingual LLMs, Teuken 7B-base and Teuken 7B-instruct, designed to embrace Europe’s linguistic diversity by supporting all 24 official languages of the European Union. Trained on a dataset comprising around 60% non-English data and utilizing a custom multilingual tokenizer, our models address the limitations of existing Large Language Models that predominantly focus on English or a few high-resource languages. We detail the models’ development principles, i.e., data composition, tokenizer optimization, and training methodologies. The models demonstrate strong performance across multilingual benchmarks, as evidenced by their performance on European versions of ARC, HellaSwag, and TruthfulQA.
[…]
123
@@ 1 Introduction
LLMs represents a disruptive technology that has the potential to be applied in numerous applications. To develop LLMs, expertise in various areas is required, starting from large-scale data pre-processing [20], training efficient tokenizers [7, 32], pre-training the models efficiently on a vast infrastructure spanning thousands of GPUs, instruction tuning the pre-trained models, and properly evaluating them on various downstream tasks [4]. Multilingualism as an additional dimension introduces additional considerations to all phases of the model development, such as multilingual data composition and multilingual tokenizer. Therefore, it is crucial that the technology and expertise to build these models is democratized to enable different communities and organizations to employ these models for their use cases.
Many efforts in developing open-source LLMs have been undertaken, such as BLOOM [54], LLaMA-3 [4], OLMo [15], Aya [10], and Mistral [8].
While these existing efforts represent significant contributions to the community and the advancement of artificial intelligence, there are still two major limitations. First, the current open-source models are predominantly English-centric, limiting their use to a broad multilingual context covering high- and low-resource languages such as within the European Union. Relying on English-centric models that employ an English-centric tokenizer in a multilingual context introduces severe disadvantages, such as lower downstream performance, additional inference costs, and increased training costs for language-specific continued pre-training and fine-tuning for languages besides English [7, 32]. Second, open-source efforts disclose different levels of granularity in sharing details about the development of the models. Although information on the model architecture is usually shared, the decisions/ablation experiments behind certain architectural choices are not always described, which can provide crucial insights relevant to developing custom LLMs. This limitation is even more apparent in the description of the dataset composition, which often provides a coarse overview, hampering the reproduction of the work.
@@
  • We present Teuken-7B-Base [^1] https://huggingface.co/openGPT-X/Teuken-7B-base-v0.6, a European pre-trained LLM that has been trained from scratch for 6 trillion tokens based on all 24 official European languages.
  • We present Teuken-7B-Instruct [^2] https://huggingface.co/openGPT-X/Teuken-7B-instruct-v0.6, the instruction-tuned model on top of the base model to enhance instruction-following capabilities.
In addition to these artefacts, we describe our design decisions in detail to facilitate the reproduction of our work and the development of novel models based on it. Therefore, our contributions present an essential step towards the democratization of this technology across Europe.
@@ 2 Related Work
Since the introduction of GPT-3 [53], several open-source/open-weights efforts have been undertaken to train LLMs. While the large majority of the work focus on English-centric models [52, 21, 22, 15, 4], there have been also efforts training multilingual models.
One of the most prominent examples is BLOOM [54], a 176B LLM trained on 46 natural languages. Further examples that specifically address multilingualism are the encoder-decoder models mT5 [31], XLM [2], XLM-R [3], and the encoder model mBERT [24].
Unlike the previously mentioned efforts, we specifically address 24 official European languages and ensure that a significant fraction of the training data comprises non-English data, representing a major step towards European LLMs. Concurrent to our work, EuroLLM [64], a 1.7B and 9B decoder-only LLM that follows the same spirit as our undertaking by addressing all 24 European languages and Salamandra [^3] https://huggingface.co/BSC-LT/salamandra-7b covering 32 languages, have been presented.
@@ 3 Pre-Training Data
We used the dataset presented in [34] as a base since it i.) covers all 24 official EU languages, ii.) is large, iii.) contains a large amount of non-English data, iv.) comprises various domains, and v.) is filtered based on established practices. In the following, we describe the composition of our final training comprised of web-crawled and curated data.
For the web crawled part, we sampled our dataset based on the 60 filtered Common Crawls WET dumps (cf.  Section˜A.1) provided by Brandizzi et al. [34]. The raw WET dumps have been filtered based on established heuristics, ensuring that the pre-training corpus is of high quality [34]. Additionally, we employed the recently released Fine-Web EDU [20] and DCLM [26] dataset since significant performance improvements for models trained on these datasets have been reported. It should be noted that Fine-Web EDU is only available in English. We still integrated it and investigated the cross-lingual performance of the corresponding model checkpoint (see Section˜5.2). To ensure a multilingual composition, we up-sampled all languages except English, which we down-sampled.
In addition to web crawled data, we included all curated datasets from [34]. Part of the curated datasets are domain-specific data such as academics, finance, and patents ensuring the diversity of our training dataset.
Our composed training dataset contains 6 trillion tokens, of which 86.79% (cf. Table˜40) originates from web data, and the remaining 13.21% represent is curated data. As illustrated in Figure˜1 and Figure˜2, 41.70% of the tokens stem from English content and due to the inclusion of German, French, and Spanish, we approach around two-thirds of the total tokens.
@@ 4 Multilingual Tokenization and Fertility Impact on Model Efficiency
In multilingual natural language processing (NLP), it is crucial to train balanced multilingual tokenizers [7, 32] to avoid increased training and inference costs and latency during inference for non-English queries. Furthermore, it prevents the model from learning long-range dependencies in limited context windows [11].
Therefore, we developed a custom multilingual tokenizer, closely following et al. [32], that is optimized for all 24 official European languages. The tokenizer training dataset contains an equal number of documents for each of the 24 languages. The documents were sourced from the same data as used in pretraining. It aims to reduce excessive text fragmentation, a phenomenon termed high “fertility”, and refers to the average number of tokens generated per word.
Fertility ($F$) is defined as the ratio of the total number of tokens ($T$) to the total number of words ($W$) in a text, i.e., $F=\frac{T}{W}$.
We conducted a fertility analysis on 2,000 sentences from the FLORES-200 [33] dataset to compare tokenizers. Because the dataset is translated across languages, i.e., the analysis is conducted on semantic equivalent content, it provides a reliable basis for evaluation. A comparison with other widely-used tokenizers is presented in Figure˜3
@@ 5 Base Model
In the following, we describe the model architecture (Section˜5.1) and training (Section˜5.2). Additionally, we describe in the appendix the used training framework (Section˜A.6) and the training infrastructure (Section˜A.7).
@@ 5.1 Model Architecture
Our model is a 7B transformer-based decoder-only model. Table˜6 provides an overview of our model architecture. We want to highlight that our architectural choices are derived from internal ablation studies and findings from related work. Our models have a sequence length of 4096 tokens and employ Rotary [27] positional embeddings that are employed to train state-of-the-art models [4]. To accelerate inference and reduce memory requirements, we employed grouped-query attention [23]. An entire overview of our architectural choices is presented in Table˜6.
@@
These and other design decisions were guided by medium-scale (Chinchilla-optimal [25] training of a 2.6B parameter model) ablation runs for various training-related hyperparameters. Our goal with these ablations was to find improvements in the compute-equivalent setting while also confirming whether proposed modifications transfer to our codebase, which is not necessarily the case [48]. Appendix A.2 describes the conducted ablation experiments in detail, Table 5 summarizes the results, and Figures 7 and 14 contain loss curves for the various experiments.
5.2 Initial Training
@@
Using the causal language modelling training objective, we trained our model on 6T tokens covering all 24 European languages as described in Section˜3, employed AdamW as an optimizer and used a cosine learning rate schedule starting with a learning rate of 3e-5, increasing it to the maximum learning rate of 3e-4 within the first 10,000 steps, and decaying it afterwards. During the training, we took two additional design decisions motivated by recent findings from related work.
5.3 Adjusted Data Mixture
@@
Penedo et al. [20] showed that training based on educational content further improves the performance of LLMs. Therefore, after 2.85T tokens, we conducted an ablation where we trained one model up to 3T tokens based on FineWeb-EDU [20] (which is only available in English), and a second model with the initial data composition (see Section˜3). As shown in Figure˜5, the model trained based on education content obtained an average performance improvement of 3% across languages and benchmarks. In the appendix, we show the performance development for English (Figure˜15), French (Figure˜16), German (Figure˜17), Finish (Figure˜18) and Estonian (Figure˜19).
@@
Figure 5: Downstream performance of the base model from 0.99T to 6T tokens across 21 European languages. The grey area highlights the ablation comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
5.4 Continued Pre-Training
We want to emphasize that the initial goal has been to train our base model up to 3T tokens, which are beyond compute-optimal based on the scaling laws presented by Hoffmann et al. [25] for our model size. Recent works show that it is beneficial to further the pre-train model, e.g., LLaMA 3 has been trained up to 15T tokens, we decided to train the model on an additional 3T tokens and therefore needed to define a new learning rate schedule. We ablated three different learning rate schedules Full-Rewarmup, Partial-Rewarmup, and Continue Decay. In the Full-Rewarmup setting, we defined a cosine learning rate schedule with the same minimal and maximal learning as for the initial pre-training phase. In the Partial-Rewarmup, we set the maximum learning rate to a quarter of the maximum learning rate of the initial pre-training phase. Finally, in the Continue-Decay setting, we decreased the learning rate to 5% of the maximum learning rate within the following 3T tokens, providing the lowest training loss (Figure˜4). Therefore, we decided to employ this setting to train our model further.
@@ 6 Instruction Tuned Model
In the following, we describe our post-training procedure to instruction-tune Teuken-7B-Base resulting in Teuken-7B-Instruct, starting with a description of the respective fine-tuning stages (Section˜6.1), followed by details about the used dataset composition (Section˜6.2).
6.1 Post-Training
For instruction-tuning, we follow a common two-step approach with a first supervised finetuning (SFT) stage to enable the model to follow general instructions and a subsequent stage of preference optimization to refine the model’s response behaviour.
Supervised finetuning (SFT).
In SFT, models are provided with an instruction as an input and are optimized to produce a respective response as an output. Our SFT setup follows common practice, with hyperparameter settings optimized via a small grid search (for details on hyperparameters we refer to Appendix A.5.1).
In addition to the general setup, we employ noisy embeddings [35] with $\alpha$ = 5; we mask the gradient of the input sequence and only update the learning signal from the response. Ultimately, we employ sequence packing to improve fine-tuning efficiency and optimize GPU usage [51].
Preference optimization.
After the SFT stage, we further improve the model with direct-preference optimization (DPO) [43]. Again, we optimize hyperparameters on a small grid search (detail in Appendix A.5.1), and equivalent to SFT, we mask the input sequences.
@@ 6.2 Data
Our dataset comprises publicly available datasets and a part that we synthesized. Inspired by the findings of [1], we utilized multilingual datasets to improve the cross-lingual performance of our model. A list of datasets employed alongside their sample size and languages is presented in Table 38 for SFT and 39 for DPO, respectively.
Synthesized Data
For generating SFT data, we follow the self-instruct paradigm [59] with modified seed prompts and a different generation model [9] ("Sigma"). To further enhance instruction diversity and complexity, we apply Evol-Instruct [14] using Mixtral, and regenerate the responses accordingly ("Sigma Evolved"). Additionally, we generate a small set of Everyday Conversations [62] using Mixtral (“Sigma Everyday conversations”). Finally, we manually curate 113 self-awareness data points designed to inform the model about its own identity ("Teuken Self-Awareness SFT"). These are oversampled by a factor of 10 during training.
Translation
We use Mixtral-8x22B-Instruct to generate high-quality translations of various source datasets into German, French, and Italian. Mixtral was selected after evaluating multiple LLMs of different scales on a representative subset of the SFT data, where it achieved the best translation quality among models with permissive licenses. Translation quality was assessed using Llama-3.1-70B-Instruct [4] as an automated judge, showing strong agreement with an independent human annotator.
Data selection
We use the deita quality and complexity scorer [56] to evaluate all instruction–response pairs. After normalizing scores across the full dataset, we compute a general preference score for each pair as a weighted sum of the two components [^4] Weights used: $w_{\text{quality}}=0.7$, $w_{\text{complex}}=0.3$. We then sample subsets of the source datasets by prioritizing data points with high preference scores.
For datasets we consider overly homogeneous, we further filter by selecting only data points that exhibit a minimal embedding distance (based on cosine similarity) to previously chosen samples (see Min Distance in Table 38). Embeddings are obtained by encoding all instructions using a SentenceTransformer model [^5] https://hf.co/sentence-transformers/distiluse-base-multilingual-cased-v1 [36, 37].